OpenStudy (veda):

Confirm that f and g are inverses by showing that f(g(x)) = x and g(f(x)) = x. Any help would be great

OpenStudy (veda):

f(x)=4/x and g(x)=4/x

OpenStudy (jerry45):

this isnt the math section